OpenStudy (anonymous):

what is the sum of 8 / x-2 + 2 / x+3 been trying to figure this out for a while

OpenStudy (anonymous):

anyone can explain ?

OpenStudy (anonymous):

Are you just looking to simplify it?

OpenStudy (anonymous):

ya trying to find the sum

OpenStudy (anonymous):

Depends on where you put the parentheses.

OpenStudy (anonymous):

Is it (8 / x-2) + (2 / x+3), meaning it's adding two rational expressions?

OpenStudy (anonymous):

Because if it is, then you could solve it like this: 1. Create an LCD between the fractions(multiplying each fraction by the other fraction's denominator written in both the top and bottom the fraction, so it doesn't change the value) (8x+24)/(x-2)(x+3)+(2x-4)/(x-2)(x+3) 2.Add the numerators, which is possible now with the common numerator. (10x+20)/(x-2)(x+3) 3.Factor the numerator(unless otherwise specified). 10(x+2)/(x-2)(x+3)

OpenStudy (anonymous):

\[\frac{ 8 }{ x } -2+\frac{ 2 }{ x }+3 = \frac{ -2x+8 }{ x }+\frac{ 2 }{ x }+3 =\frac{ -2x+10 }{ x }+3=\frac{ x+10 }{ x }\] That's the step by step solution for exactly what you typed above. If it has parentheses, fix it and I'll try to solve that one. So, my final answer is: \[\frac{ x+10 }{ x }\]

OpenStudy (anonymous):

Ah, there needs to be more parentheses. (8 / (x-2)) + (2 / (x+3)), so that it divides it properly. I got the same problem, actually. I just forgot to note it. Sorry.

OpenStudy (anonymous):

It's all about the parentheses. What did the problem look like in it's original state?